OpenStudy (anonymous):

write an equation for the situation: Line that is perpendicular to y=-4/3x+8 and goes through the point (4,5)

OpenStudy (anonymous):

find the slop (m )

OpenStudy (anonymous):

i got y=3/4x+2... is this right?

OpenStudy (anonymous):

slop = -4/3 line is perpendicular m= 3/4 y=mx +b replace m , x, y to the formula we got : 5 = 3/4(4) +b find b

OpenStudy (anonymous):

you got right

OpenStudy (anonymous):

ok thanks